# Individualizing Glioma Radiotherapy Planning by Optimization of a Data and Physics-Informed Discrete Loss

### v3

@@ Abstract
Brain tumor growth is unique to each glioma patient and extends beyond what is visible in imaging scans, infiltrating surrounding brain tissue. Understanding these hidden patient-specific progressions is essential for effective therapies. Current treatment plans for brain tumors, such as radiotherapy, typically involve delineating a uniform margin around the visible tumor on pre-treatment scans to target this invisible tumor growth. This ”one size fits all” approach is derived from population studies and often fails to account for the nuances of individual patient conditions. We present the GliODIL framework, which infers the full spatial distribution of tumor cell concentration from available multi-modal imaging, leveraging a Fisher-Kolmogorov type physics model to describe tumor growth. This is achieved through the newly introduced method of Optimizing the Discrete Loss (ODIL), where both data and physics-based constraints are softly assimilated into the solution. Our test dataset comprises 152 glioblastoma patients with pre-treatment imaging and post-treatment follow-ups for tumor recurrence monitoring. By blending data-driven techniques with physics-based constraints, GliODIL enhances recurrence prediction in radiotherapy planning, challenging traditional uniform margins and strict adherence to the Fisher-Kolmogorov partial differential equation (PDE) model, which is adapted for complex cases.
@@ 1 Main
Gliomas are the most common primary brain tumors in adults. [1, 2]. Commonly used treatment strategies include surgery, chemotherapy, and radiotherapy. Despite advances in understanding the biological basis of these tumors and the multi-modal combination of therapies, the prognosis of glioma patients, in particular those with glioblastoma (WHO-CNS grade 4) [3], remains dismal. A key challenge for more successful therapy of glioma patients is the infiltrative tumor growth pattern. Already at initial diagnosis, glioma cells have invaded the surrounding brain parenchyma well beyond the tumor margins visible on conventional imaging. To account for this otherwise invisible tumor growth, both North American and European guidelines for radiotherapy planning define standard, uniform safety margins around the resection cavity and/or remaining tumor on conventional Magnetic Resonance Imaging (MRI), a technique that utilizes magnetic fields and radio waves to generate detailed images of the internal structures of the body [4, 5]. Despite many efforts, truly tailoring radiotherapy to an individual patient’s tumor’s spread is an unmet clinical need in neurooncology [6].
@@
Computational modeling has potential to improve radiotherapy volume definition [7, 8], particularly through the use of Partial Differential Equations (PDEs) to simulate the nuanced progression of tumors. This approach, exemplified by models like the Fisher-Kolmogorov (FK) equation, enables the prediction of spatial tumor cell distribution by accounting for migration and proliferation. While more intricate diffusion-reaction models introduce the complexity of various tumor cell states, such as necrotic and proliferative, they also demand extensive data for precise parameter estimatiom. Incorporating multimodal imaging, including Fluoroethyl-L-Tyrosine Positron Emission Tomography (FET-PET), our framework aims to refine predictions of tumor infiltration patterns. FET-PET provides metabolic information that, when combined with MRI’s structural data, offers a comprehensive view of tumor behavior. This integration not only parallels dose boosting strategies in targeting areas of high metabolic activity but also underscores the potential of using such detailed imaging data to inform models that predict tumor growth and spread with greater accuracy.
@@
Existing approaches to personalizing tumor growth models require solving the inverse problem, i.e., inferring the growth model parameters providing an optimal fit to the clinically observed tumor on imaging [9, 10, 11, 12, 13, 14]. Given that we only have a single time point of data when initiating treatment, the growth models employed must be exceedingly simple. Without this simplicity, the problem becomes highly ill-posed, and accurate retrieval of parameters becomes unattainable. In addition, traditional methods, e.g., those based on Monte Carlo sampling [15], have a severe limitation, namely, long computational time to perform parametric inference. Both the computational time and simplistic tumor growth equations severely limit clinical applicability of such models for radiotherapy planning. Recently, deep learning methods were introduced to address the computational time issue[16, 17, 18, 19, 20, 21, 22]. While these learnable methods offer improvements in computational efficiency, their current lack of robust error control and solution accuracy poses significant challenges in medical applications, where the utmost precision is required for treatments and patient care. Until these issues are adequately addressed, the use of such models in critical medical decisions remains limited.
Among the fastest traditional methods, PDE-constrained optimization has been extensively applied for tumor growth calibration. This approach leverages Partial Differential Equations (PDEs) to impose constraints within an optimization framework, enabling the modeling of complex physical phenomena such as tumor progression. Notable efforts, such as the work by Subramanian et al.[14], utilize biophysically motivated regularization to determine initial tumor conditions and estimate proliferation and infiltration rates. This method involves running multiple simulations with varying parameters until finding a set that adequately fits the observed data. However, this technique is firmly anchored to a predefined PDE model and its assumptions regarding initial conditions, leading to a rigidity that may not fully capture the intricate dynamics of tumor biology. As a result, these models often struggle to accommodate discrepancies between formal model assumptions and the actual patterns of tumor growth. This scenario highlights the need for a methodology that melds model-driven and data-driven approaches, aiming for a balanced compromise.
@@
Physics-Informed Neural Networks (PINNs) emerge as a middle ground, aiming to strike a balance between the rigidity of PDE models, initial conditions, and the flexibility of data-driven approaches. They embed physical laws in the form of differential equations directly into the architecture of neural networks [23]. This theoretically allows for more reliable and physically meaningful predictions by using the neural network to approximate the solution to the PDE and adapt the initial condition accordingly to soft imposed assumptions. However, the practical application of PINNs in a clinical setting is currently challenging due to computational inefficiencies. Modifying a single weight in these often densely connected networks can have a widespread, non-local impact on their output, making calibration a notoriously difficult task. Moreover, although the PDE residual serves as a penalty term in PINNs, its testing is confined to a restricted number of points rather than being evaluated globally across the entire computational domain. Hence, while PINNs offer a promising avenue for model personalization, they currently fall short in terms of computational feasibility for real-world applications.
@@
In summary, three key challenges must be addressed to facilitate the successful clinical translation of computational tumor growth models: (i) enhancing computational efficiency to enable timely and practical use in a clinical setting, (ii) establishing robust error control to ensure safe clinical application, and (iii) introducing the flexibility to adapt models beyond basic mathematical frameworks, thereby capturing the complexities of tumor growth more accurately. This necessitates a balance between adhering to the growth model and accurately reflecting the tumor observed in clinical practice.
Figure 1: GliODIL overview. a Multi-modal patient data consisting of MR and PET imaging. b To define the input grid conditions and tumor boundaries, tissue extraction (using an atlas registration; left) [15] and automatic brain tumor segmentation (right) [24] is performed. Automatically segmented tumor regions include three components:(i) edema, characterized by tissue swelling due to fluid accumulation; (ii) enhancing core, indicative of active tumor growth and characterized by vascular leakage, and (iii) necrotic core, showing tissue death due to hypoxia or nutrient deprivation. Corresponding FET-PET scans provide metabolic insight, further aiding in accurate tumor delineation. c,d Spatio-temporal progression of a tumor within patient anatomy. Values are stored on a 4-dimensional multi-resolution grid. Optimization utilizes automatic differentiation of each gridpoint and is guided by the loss function. Since we model the progression based on a single time-point input data, the growth parameters are being resolved up to a timescale. Calculation of discrepancy between patient’s tumor characteristics $L_{\text{DATA}}$ and proposed by GliODIL tumor cell-distribution’s final time-point. Tumor growth physics is encapsulated in a partial differential equation (PDE). Calculation of PDE’s residual $L_{\text{PDE}}$ and single focal initial condition $L_{\text{IC}}$. e GliODIL outputs. The framework successfully infers the complete distribution of tumor cells, facilitating the development of a radiotherapy plan. This plan effectively covers areas of tumor recurrence identified in post-operative data, while maintaining the total radiotherapy volume in line with standard clinical guidelines.
@@
This work introduces Glioma ODIL (GliODIL), a novel optimization framework designed for estimating tumor cell concentrations and migration pathways surrounding visible tumors in MRI and PET scans, as depicted in Figure 1. Our approach uniquely integrates traditional numerical methods with data-driven paradigms, providing a more comprehensive insight into tumor behavior. In addition to its adaptive capabilities. GliODIL builds upon our previous work on the Optimizing a Discrete Loss (ODIL) technique [25, 26] which significantly enhances computational speed compared to traditional PINN architectures. By utilizing consistent and stable discrete approximations of the PDEs, employing a multi-resolution grid, and leveraging automatic differentiation, we achieve computation times suitable for clinical applications such as radiotherapy planning.
@@
Diverging from conventional glioma models, which primarily adjust parameters within predetermined PDE frameworks, our GliODIL methodology uniquely refines both the parameters and the discretized solutions. This refinement process involves optimizing a cost function that seamlessly integrates the growth equations in their discrete form with empirical data, treating these elements as soft constraints. This approach not only enhances computational efficiency by streamlining the optimization process but also incorporates robust error control mechanisms. These mechanisms assess the model’s fidelity by calculating the physics residual error and evaluating how closely our assumptions align with the tumor progression reflected in the clinical data. Furthermore, GliODIL introduces the necessary flexibility to move beyond basic mathematical models, thereby more accurately capturing the intricate realities of tumor growth and progression.
@@
To substantiate these advancements, we present dedicated experiments on clinical data, demonstrating GliODIL’s capacity to address the outlined challenges effectively. Our test dataset encompasses 152 glioblastoma patients, including 58 cases with pre-treatment FET-PET imaging alongside MRI, and post-treatment MRI follow-ups for comprehensive tumor recurrence monitoring and model validation. These follow-ups occur within a timeframe typically ranging from 1 month to 1 year after treatment, providing a broad spectrum for assessing GliODIL’s performance in real-world clinical scenarios. By making this dataset publicly available, we not only validate GliODIL’s efficacy in enhancing tumor recurrence coverage but also contribute to the broader pursuit of personalized treatment strategies in oncology. This initiative challenges the constraints of rigid PDE models and the traditional reliance on uniform margins.
@@ 2 Results
Our study initiates by inferring full spatial tumor cell distributions, initially employing a synthetic dataset to evaluate our model’s performance under varied tumor conditions and to fine-tune its parameters. This involves examining cases of individual tumors and localized multi-focal tumors, demonstrating the model’s capacity to handle noise within the synthetic data and scenarios that resemble post-resection conditions. Subsequently, the model is applied to data collected from 152 real patients (discussed in Section 4.6) to verify its capacity for accurately inferring tumor growth parameters that depict the patient-specific scenarios, and benchmarks these findings against established methodologies.
In addressing the inverse problem of glioma modeling, we compare our results with those obtained from the Covariance Matrix Adaptation Evolution Strategy (CMA-ES) method, detailed in [27, 28], which relies on numerous simulations to identify parameters that best fit the data, and the Learn-Morph-Infer (LMI) technique [22], employing a deep learning framework for tumor growth model parameter estimation. For clarity, solutions using the forward PDE finite difference method or finite element method with parameters obtained through CMA-ES and LMI are labeled as $\text{PDE}_{\text{CMA-ES}}$ and $\text{PDE}_{\text{LMI}}$, respectively, emphasizing that they directly solve the PDE.
@@
Further assessment of our model involves its application in creating personalized radiotherapy plans. These plans are evaluated against commonly used uniform margin radiation therapy plans and other prevalent modeling techniques to assess the clinical utility of our approach. Our analysis integrates cell concentrations directly measured from GliODIL and the outcomes from forward PDE simulations using parameters estimated by GliODIL, denoted as $\text{PDE}_{\text{GliODIL}}$. This indicates a strict adherence to the Tumor Growth Model’s PDE. The differences in these methodologies, especially in the context of complex tumors, are thoroughly examined in 4. To enhance the efficiency of GliODIL, we implement a personalized initial guess for the tumor cell distribution, further elaborated upon in Section 4.7.
2.1 Full Spacial Tumor Cell Distribution Inference
@@
Our main modeling assumption is that tumor growth according to a given PDE is controlled by the loss $L_{\text{PDE}}$ and that it starts from a single focal seed point, controlled by the initial condition $L_{\text{IC}}$ loss. Both of these loses are formally introduced in Section 4.1. We want to test the applicability of these assumptions to more complex, real-world scenarios like localized multi-focal tumors that break the modeling assumptions. We considered relaxing both assumptions to capture such complex scenarios. As the initial condition of the tumor is uncertain and may not necessarily begin from a single focal point, we assign a low importance to it in our GliODIL solution. Further easing the constraints imposed by $L_{\text{IC}}$ results in ambiguities that hinder the reliable determination of tumor growth parameters of single focal tumors, a core aspect of our study. Instead, we opted for experimenting with the $L_{\text{PDE}}$ importance governed by a weighing parameter $\lambda_{\text{PDE}}$ in the final loss.
@@
We report the results from 100 synthetic patients, half with single focal tumors and the remaining half with multi-focal. The outcomes of this experiment are illustrated in Figure 2. In Figure 2e we observe that both high $\lambda_{\text{PDE}}$ (strong adherence to the equation) and low $\lambda_{\text{PDE}}$ (overfitting to the data) are sub-optimal for complex tumors and the $\lambda_{\text{PDE}}=1$ used in GliODIL aligns the closest with the ground truth. The decision to set $\lambda_{\text{PDE}}=1$ represents a balanced choice, enabling accurate inference of single focal tumors (see Figure 2a,b,c) while also capturing a significant portion of the multi-focal ground truth, as shown in Figure 2a,e. This setting of $\lambda_{\text{PDE}}=1$ will be consistently applied in our subsequent experiments involving real patient data.
Figure 2: Synthetic dataset experiments overview: calibration of PDE loss importance. a Illustration of GliODIL’s synthetic results for localized multi-focal and single focal tumors. In these simulations, all loss weights remain constant except for the $L_{\text{PDE}}$ weight, $\lambda_{\text{PDE}}$. A value of $\lambda_{\text{PDE}}=1$ was selected to strike a balance in the model between fit to the data and the tumor growth equations. Dice scores were calculated at 10% tumor cell concentration, which is outside of the segmented region (edema at least at 20%). b The first and second rows represent a segmentation and a synthetic FET-PET that serve as input to GliODIL. Comparison: a forward run with ground truth parameters (Reference), tumor cells distribution inferred by GliODIL, and an initial guess (I. guess) that serves as a starting point of the optimization. c Comparison of exact and inferred parameters. d Input for the multi-focal experiment. e Tumor cell distribution of GliODIL with progressively relaxed $\lambda_{\text{PDE}}$. Root Mean Square Error (RMSE) is calculated for low-cell tumor concentration (at 10%), meaning tumor cell distribution outside of visible tumor on MRI.
@@
Regarding data explanation, as measured by dice scores and PET signal correlation, GliODIL surpasses all examined PDE solutions. Better fit to the data can be attributed to GliODIL’s ability to balance between adhering to the PDE on which it is being regularized on and effectively explaining the data. This performance indicates that its inferred tumor cell distributions could more accurately mirror actual conditions. However, for real patients, unlike in synthetic experiments detailed in Section 2.1, we lack ground truth to directly substantiate these claims, and the results might be overfitted through the data-driven term (as in Figure 2e for $\lambda_{\text{PDE}}<1$). To validate GliODIL’s performance, a downstream task of clinical relevance is performed. In the following Section 2.2, we demonstrate that GliODIL leads to more effective radiotherapy plans regarding tumor recurrence coverage.
Figure 3: Real patient data analysis: tumor cell concentration inference.. a, b Comparison of tumor cell density predictions from various models with corresponding data inputs. Threshold segmentation values for Dice scores for each model are determined through a grid search, since the LMI model does not infer them. c Average data-fit scores for each model.
2.2 Radiotherapy Planning and Recurrence coverage

### v4

@@ Abstract
Brain tumor growth is unique to each glioma patient and extends beyond what is visible in imaging scans, infiltrating surrounding brain tissue. Understanding these hidden patient-specific progressions is essential for effective therapies. Current treatment plans for brain tumors, such as radiotherapy, typically involve delineating a uniform margin around the visible tumor on pre-treatment scans to target this invisible tumor growth. This ”one size fits all” approach is derived from population studies and often fails to account for the nuances of individual patient conditions. We present the GliODIL framework, which infers the full spatial distribution of tumor cell concentration from available multi-modal imaging, leveraging a Fisher-Kolmogorov type physics model to describe tumor growth. This is achieved through the newly introduced method of Optimizing the Discrete Loss, where both data and physics-based constraints are softly assimilated into the solution. Our test dataset comprises 152 glioblastoma patients with pre-treatment imaging and post-treatment follow-ups for tumor recurrence monitoring. By blending data-driven techniques with physics-based constraints, GliODIL enhances recurrence prediction in radiotherapy planning, challenging traditional uniform margins and strict adherence to the Fisher-Kolmogorov partial differential equation model, which is adapted for complex cases.
@@ 1 Introduction
Gliomas are the most common primary brain tumors in adults. [vandenBent2023, Ostrom2023]. Commonly used treatment strategies include surgery, chemotherapy, and radiotherapy. Despite advances in understanding the biological basis of these tumors and the multi-modal combination of therapies, the prognosis of glioma patients, in particular those with glioblastoma (WHO-CNS grade 4) [Louis2021], remains dismal. A key challenge for more successful therapy of glioma patients is the infiltrative tumor growth pattern. Already at initial diagnosis, glioma cells have invaded the surrounding brain parenchyma well beyond the tumor margins visible on conventional imaging. To account for this otherwise invisible tumor growth, both North American and European guidelines for radiotherapy planning define standard, uniform safety margins around the resection cavity and/or remaining tumor based on conventional Magnetic Resonance Imaging (MRI), which highlights areas affected by edema alongside regions of actively growing and necrotic tumor tissue [Weller2021, Niyazi2023]. Despite many efforts, truly tailoring radiotherapy to an individual patient’s tumor’s spread is an unmet clinical need in neurooncology [Frosina2023].
In current clinical practice, radiotherapy planning for glioblastoma patients typically employs a uniform safety margin of 1.5 centimeters [Wen2010] around the visible tumor core identified in preoperative Magnetic Resonance Imaging (MRI) scans. This standardized approach aims to encompass both the detectable tumor and potential microscopic extensions that are not visible on imaging. However, this uniform margin does not account for the highly heterogeneous nature of tumor infiltration, variations in individual brain anatomy, or the presence of anatomical barriers that can influence tumor spread. Consequently, the uniform margin may either over-treat healthy brain tissue, leading to unnecessary side effects, or under-treat regions where the tumor has infiltrated beyond the prescribed margin, increasing the risk of early recurrence. These shortcomings highlight the urgent clinical motivation to seek more personalized and precise radiotherapy planning methods that can adapt to the unique tumor dynamics and anatomical conditions of each patient.
@@
Computational modeling has the potential to improve the definition of radiotherapy volume [cristini2010multiscale, mang2020integrated], particularly through the use of partial differential equations (PDE) to simulate the nuanced progression of tumors. This approach, exemplified by models like the Fisher-Kolmogorov (FK) equation, enables the prediction of spatial tumor cell distribution by accounting for migration and proliferation. Although more intricate diffusion-reaction models introduce the complexity of various tumor cell states, such as necrotic and proliferative, they also require extensive data for precise parameter estimation. Incorporating multimodal imaging, including Fluoroethyl-L-Tyrosine Positron Emission Tomography (FET-PET), ourframework aims to refine predictions of tumor infiltration patterns. FET-PET provides metabolic information [Schon2020, Kunz2011] that, when combined with structural data from MRI, offers a complete view of tumor behavior. This integration not only parallels dose-boosting strategies in targeting areas of high metabolic activity but also underscores the potential of using complex multimodal imaging data to inform models that predict tumor growth and spread with greater precision.
@@
Existing approaches to personalizing tumor growth models require solving the inverse problem, i.e., inferring the growth model parameters that provide an optimal fit to the clinically observed tumor on images [hogea2008image, menze2011generative, konukoglu2010image, unkelbach2014radiotherapy, le2015bayesian, knopoff2017mathematical, b6, scheufele2020image, Subramanian2020_b7]. Given that we only have a single time point of data when initiating treatment, the growth models employed must be exceedingly simple, leading to significant mismatch between model parameters and image observations, and requiring strong assumption about initial conditions. Without this simplicity, the problem becomes highly ill-posed, and accurate retrieval of parameters becomes unattainable. In addition, traditional methods, e.g., those based on Monte Carlo sampling [lipkova2019personalized], have a severe limitation, namely, long computational time to perform parametric inference. Both the computational time and simplistic tumor growth equations severely limit clinical applicability of such models for radiotherapy planning.
Recently, deep learning methods were introduced to address the computational time issue of tumor cells inference [petersen2019deep, petersen2021continuous, mascheroni2021improving, wang2022biologically, nardini2020learning, ezhov2021geometry, ezhov2022learn]. While these learnable methods offer improvements in computational efficiency, their current lack of robust error control and solution accuracy poses significant challenges in medical applications, where the utmost precision is required for treatments and patient care. Until these issues are adequately addressed, the use of such models in critical medical decisions remains limited.
@@
Physics-Informed Neural Networks (PINNs) emerge as a middle ground, aiming to strike a balance between the rigidity of PDE models, initial conditions of the tumor cells spatial distribution, and the flexibility of data-driven approaches. They embed physical laws in the form of differential equations directly into the architecture of neural networks [raissi2019physics]. Theoretically, this allows for more reliable and physically meaningful predictions by using the neural network to approximate the solution to the PDE and adapt the initial condition accordingly to soft imposed assumptions. However, the practical application of PINNs in a clinical setting is currently challenging due to computational inefficiencies. Modifying a single weight in these often densely connected networks can have a widespread, non-local impact on their output, making calibration a notoriously difficult task, limiting its clinical applicability. Moreover, although the PDE residual serves as a penalty term in PINNs, its testing is confined to a restricted number of points rather than being evaluated globally across the entire computational domain. Hence, while PINNs offer a promising avenue for model personalization, they currently fall short in terms of computational feasibility for real-world applications.
@@
In summary, three key challenges must be addressed to facilitate the successful clinical translation of computational tumor growth models: (i) enhancing computational efficiency to enable timely and practical use in a clinical setting, (ii) establishing error control to ensure safe clinical application, and (iii) introducing the flexibility to adapt models beyond basic mathematical frameworks, thereby capturing the complexities of tumor growth more accurately. This necessitates a balance between adhering to the growth model and accurately reflecting the tumor observed in clinical practice.
@@
This work introduces Glioma ODIL (GliODIL), a novel optimization framework designed for estimating tumor cell concentrations and migration pathways surrounding visible tumors in MRI and PET scans, as depicted in Figure 1. Our approach uniquely integrates traditional numerical methods with data-driven paradigms, providing a more comprehensive insight into tumor behavior. In addition to its adaptive capabilities. GliODIL builds upon our previous work on the Optimizing a Discrete Loss (ODIL) technique [karnakov2024odil, Karnakov2023] which significantly enhances computational speed compared to traditional PINN architectures. By utilizing consistent and stable discrete approximations of the PDEs, employing a multi-resolution grid, and leveraging automatic differentiation, we achieve computation times suitable for clinical applications such as radiotherapy planning.
@@
Diverging from conventional glioma models, which primarily adjust parameters within predetermined PDE frameworks, our GliODIL methodology uniquely refines both the parameters and the discretized solutions. This refinement process involves optimizing a cost function that seamlessly integrates the growth equations in their discrete form with empirical data, treating these elements as soft constraints. This approach not only enhances computational efficiency by streamlining the optimization process but also incorporates error control mechanisms. These mechanisms assess the model’s fidelity by calculating the physics residual error and evaluating how closely our assumptions align with the tumor progression reflected in the clinical data. Furthermore, GliODIL introduces the necessary flexibility to move beyond basic mathematical models, thereby more accurately capturing the intricate realities of tumor growth and progression.
@@
We present dedicated experiments on clinical data, demonstrating GliODIL’s capacity to address the outlined challenges effectively. Our test dataset encompasses 152 glioblastoma patients, including 58 cases with pre-treatment FET-PET imaging alongside MRI, and post-treatment MRI follow-ups for comprehensive tumor recurrence monitoring and model validation. These follow-ups occur within a timeframe typically ranging from 1 month to 1 year after treatment, providing a broad spectrum for assessing GliODIL’s performance in real-world clinical scenarios. By making this dataset publicly available, we not only validate GliODIL’s efficacy in enhancing tumor recurrence coverage but also contribute to the broader pursuit of personalized treatment strategies in oncology. This initiative challenges the constraints of rigid PDE models and the traditional reliance on uniform margins.
Figure 1: GliODIL Overview. a Multi-modal patient data comprising MR and PET imaging. Tissue extraction using atlas registration [lipkova2019personalized] and automatic brain tumor segmentation [kofler2020brats] are performed to define the tumor boundaries and microenvironments. Automatically segmented tumor regions include three components:(i) edema, characterized by tissue swelling due to fluid accumulation; (ii) enhancing core, indicative of active tumor growth and characterized by vascular leakage, and (iii) necrotic core, showing tissue death due to hypoxia or nutrient deprivation. Corresponding FET-PET scans provide metabolic insight, further aiding in accurate tumor delineation. b Prior knowledge about the tumor growth process and the imaging signatures of tumor cells. The physics of tumor growth is described by a partial differential equation (PDE), while the relationship between tumor cells and the available imaging data is modeled through the Imaging Model. c Spatio-temporal progression of a tumor within patient anatomy. Calculation of PDE’s residual $L_{\text{PDE}}$ and single focal initial condition $L_{\text{IC}}$. Unknown fields are stored on a 4-dimensional multi-resolution grid. Optimization utilizes automatic differentiation of each gridpoint and is guided by the loss function. Since we model the progression based on a single time-point input data, the growth parameters are being resolved up to a timescale. Calculation of discrepancy between patient’s tumor characteristics $L_{\text{DATA}}$ and proposed by GliODIL tumor cell-distribution at the final time-point. d GliODIL outputs. The framework successfully infers the complete distribution of tumor cells, facilitating the development of a radiotherapy plan. This plan effectively covers areas of tumor recurrence identified in post-operative data, while maintaining the total radiotherapy volume in line with standard clinical guidelines.
@@ 2 Results
To provide context for our experimental findings, we begin with a concise overview of the GliODIL pipeline (Figure 1), illustrating how input data, prior knowledge of tumor growth, and PDE-based constraints combine to yield patient-specific predictions of tumor evolution and inform radiotherapy planning.
As shown in Figure 1(a), GliODIL operates on anatomical and metabolic information derived from MR and FET-PET imaging. Tissue extraction using atlas registration [lipkova2019personalized] and automatic tumor segmentation [kofler2020brats] delineate the tumor’s boundaries (edema, enhancing core, and necrotic core), providing the necessary spatial context for subsequent analysis.
Figure 1(b) highlights how physics- and imaging-based constraints are integrated. Specifically, the Tumor Growth Model (Section 4.1) uses a Fisher-Kolmogorov Reaction-Diffusion (FK) PDE to capture cellular diffusion and proliferation, and the Imaging Model (Section 4.3) links these tumor cell distributions to the observed MR and PET data.
As depicted in Figure 1(c), the model estimates the tumor’s 4D evolution—storing unknown fields in a multi-resolution grid—while minimizing a comprehensive loss function (Section 4.4). This includes enforcing PDE constraints ($L_{\text{PDE}}$), specifying focal initial conditions ($L_{\text{IC}}$), and capturing alignment with patient imaging ($L_{\text{DATA}}$). The ODIL framework (Section 4.2) guides this optimization, balancing data fidelity against physics-based regularization.
Finally, Figure 1(d) illustrates the outcome of this optimization. GliODIL infers the full tumor cell distribution within the brain and provides a basis for improved radiotherapy planning. By preserving standard clinical dose thresholds while covering areas at high risk of recurrence, GliODIL yields treatment volumes comparable to conventional margins but with better targeted coverage of complex tumor shapes.
After introducing this streamlined workflow, we evaluate GliODIL’s performance on synthetic data (Section 2)—enabling parameter fine-tuning and robust testing—and subsequently on real tumor cases from 152 patients (Section 2.1). These analyses compare GliODIL against standard practices (the “Standard Plan”) and multiple baseline models (Section 4.6), employing Dice score, FET-PET correlation, and Recurrence Coverage (Section 4.5) to demonstrate GliODIL’s predictive accuracy. GliODIL is the only model that consistently outperforms the Standard Plan.
The study is organized as follows: In Section 2, synthetic data is used to evaluate model performance under varied tumor conditions, allowing for parameter fine-tuning against a known ground truth. These conditions involve cases of individual tumors and localized multi-focal tumors, demonstrating the model’s robustness in handling noise and scenarios resembling post-resection conditions. Subsequently, in Section 2.1, GliODIL is applied to and compared against multiple baselines in real tumor cases from 152 patients (discussed in Section 4) to assess its accuracy in inferring tumor growth parameters that depict patient-specific scenarios, benchmarking these findings against established methods described in Section 4.6. For radiotherapy planning, we compare our results with the standard clinical practice of applying uniform safety margins around pre-operative MRI scans, referred to as the Standard Plan (Clinical Target Volume, CTV).
@@
Our analysis integrates cell concentrations measured directly from GliODIL and outcomes from forward PDE simulations using parameters estimated by GliODIL, denoted as $\text{PDE}_{\text{GliODIL}}$. The differences in these methodologies, especially for complex tumors, are examined in Figure 4. To enhance GliODIL’s efficiency, we implement a personalized initial guess for tumor cell distribution, detailed in Section 4.8.
Full Spatial Tumor Cell Distribution Inference
@@
Our main modeling assumption is that tumor growth according to a given PDE is controlled by the loss $L_{\text{PDE}}$ and that it starts from a single focal seed point, controlled by the initial condition $L_{\text{IC}}$ loss. Both of these loses are formally introduced in Section 4.1. We want to test the applicability of these assumptions to more complex, real-world scenarios like localized multi-focal tumors that break the modeling assumptions. We considered relaxing both assumptions to capture such complex scenarios. As the initial condition of the tumor is uncertain and may not necessarily begin from a single focal point, we assign a low importance to it in our GliODIL solution. However, we experiment with the $L_{\text{PDE}}$ importance governed by a weighing parameter $\lambda_{\text{PDE}}$ in the final loss.
@@
We report the results from 100 synthetic patients, half with single focal tumors and the remaining half with multi-focal. The outcomes of this experiment are illustrated in Figure 2. In Figure 2e we observe that both high $\lambda_{\text{PDE}}$ (strong adherence to the equation) and low $\lambda_{\text{PDE}}$ (overfitting to the data) are sub-optimal for complex tumors and the $\lambda_{\text{PDE}}=1$ used in GliODIL aligns the closest with the ground truth.
The decision to set $\lambda_{\text{PDE}}=1$ represents a balanced choice, enabling accurate inference of single focal tumors (see Figure 2a,b,c) while also capturing a significant portion of the multi-focal ground truth, as shown in Figure 2a,e. This setting of $\lambda_{\text{PDE}}=1$ will be consistently applied in our subsequent experiments involving real patient data. Notably, the balance between adherence to the data and physics priors was determined using noisy synthetic data with broken assumptions rather than patient data. This approach prevents partial dilution of the validation dataset. If additional data becomes available, some recurrence cases could potentially be used to refine the calibration of $\lambda$ values.
Figure 2: Calibration of PDE Loss Weight in Synthetic Dataset Experiments for Tumor Segmentation. a Illustration of GliODIL’s synthetic results for localized multi-focal and single focal tumors. In these simulations, all loss weights remain constant except for the $L_{\text{PDE}}$ weight, $\lambda_{\text{PDE}}$. A value of $\lambda_{\text{PDE}}=1$ was selected to strike a balance in the model between fit to the data and the tumor growth equations. Dice scores were calculated at 10% tumor cell concentration, which is outside of the segmented region (edema at least at 20%). b The first and second rows represent a segmentation and a synthetic FET-PET that serve as input to GliODIL. Comparison: a forward run with ground truth parameters, tumor cells distribution inferred by GliODIL, and an initial guess (I. guess) that serves as a starting point of the optimization. c Comparison of exact and inferred parameters. d Input for the multi-focal experiment. e Tumor cell distribution of GliODIL with progressively relaxed $\lambda_{\text{PDE}}$. Root Mean Square Error (RMSE) is calculated for low-cell tumor concentration (at 10%), meaning tumor cell distribution outside of visible tumor on MRI.
@@
Regarding data explanation, as measured by Dice scores and PET signal correlation, GliODIL surpasses all examined PDE solutions. Better fit to the data can be attributed to GliODIL’s ability to balance between adhering to the PDE on which it is being regularized on and effectively explaining the data. This performance indicates that its inferred tumor cell distributions could more accurately mirror actual conditions. However, for real patients, unlike in synthetic experiments detailed in Section 2, we lack ground truth to directly substantiate these claims, and the results might be overfitted through the data-driven term (as in Figure 2e for $\lambda_{\text{PDE}}<1$). To validate GliODIL’s performance, we conducted a downstream task with direct implications for clinical applications. In the following Section 2.1, we demonstrate that GliODIL leads to more effective radiotherapy plans regarding tumor recurrence coverage.
Figure 3: Tumor Cell Concentration Inference in Real Patient Data: Comparative Analysis of Predictive Models. a, b Comparison of tumor cell density predictions from various models with corresponding data inputs. Threshold segmentation values for Dice scores for each model are determined through a grid search, since the LMI model does not infer them. c Average data-fit scores for each model.
2.1 Clinical Dataset Validation: Radiotherapy Planning and Recurrence Coverage